OpenStudy (anonymous):

5 = –(–z + 3) Medal and fan.

OpenStudy (shreenithaa):

what do u need

OpenStudy (anonymous):

5 = –(–z + 3)

OpenStudy (shreenithaa):

the answer to x?

OpenStudy (shreenithaa):

sorry z

OpenStudy (anonymous):

Yes.

OpenStudy (haleyelizabeth2017):

First, multiply -z and 3 by -1 :)

OpenStudy (anonymous):

Don't I need Z first, though?

OpenStudy (shreenithaa):

first use the distributive property to distribute the negative sign

OpenStudy (haleyelizabeth2017):

No...distribute negative one first. What do you get when: \(-1*-z\) and \(-1*3\) ?

OpenStudy (anonymous):

Would that make it -5 + 3? (Sorry, I'm kind of bad at math.)

OpenStudy (haleyelizabeth2017):

So...two negatives equal a positive. When you multiplied -z by -1, you get z.

OpenStudy (haleyelizabeth2017):

-1*3 is...

OpenStudy (anonymous):

Uggh, I'm confused 3:

OpenStudy (haleyelizabeth2017):

lol -1*3 is -3 :) so what we have now is...: 5=z-3 Do you know how to isolate z now?

OpenStudy (anonymous):

Move it alll the way to the right?

OpenStudy (haleyelizabeth2017):

no..lol add three to both sides. The threes will cancel on the right so all you're left with is z on the right. 5=z-3 +3 +3 z=5+3

OpenStudy (haleyelizabeth2017):

Simplify and you're all set.